# Revealing Multimodal Contrastive Representation Learning through Latent Partial Causal Models

### v1

@@ 3.1 Identifiability Analysis on Hypersphere
On hypersphere, we parameterize the proposed latent partial causal generative models depicted in Figure 1 by the following:
$p(\mathbf{z}_{x})=|\mathcal{Z}|^{-1},\quad p(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{p}e^{(k\mathbf{z}^{T}_{t}\mathbf{z}_{x})},$ (4)
$\mathbf{x}=\mathbf{g}_{x}(\mathbf{z}_{x},\mathbf{m}_{x}),\quad\mathbf{t}=\mathbf{g}_{t}(\mathbf{z}_{t},\mathbf{m}_{t}),$ (5)
where $\mathcal{Z}$ denotes the space of latent factors $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$. Influenced by the commonly used feature normalization in constrastive loss, we assume that $\mathcal{Z}$ is the unit hypersphere $\mathbb{S}^{M-1}$. We do not enforce any further assumptions for $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. For $\mathbf{g}_{x}$ and $\mathbf{g}_{t}$, we assume them to be invertible (i.e., injective) mapping, ensuring the information in latent space can be recovered. In addition, we assume that $p(\mathbf{z}_{x})$ follows a uniform distribution, and $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ follows a von Mises-Fisher (vMF) distribution, considering the constraint of unit hypersphere.
Given the aforementioned assumptions, our subsequent discussion aims to establish that the minimization of the symmetric contrastive loss (as defined in Equation (2)) converges to a symmetric cross entropy, as follows:
@@
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eq. (4)-(5) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized symmetric contrastive loss given by Eq. (2) leads to the following reduction of the loss itself:
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$
$\quad\quad\quad+\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}},$ (6)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)},$ (7)
$q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})e^{(\mathbf{h}(\mathbf{z}_{t})^{T}\mathbf{h}(\mathbf{z}_{x})/\tau)},$ (8)
@@
Proof sketch
The proof can be done by the following three distinct steps, each building upon the previous one. Initially, we establish that the minimization of the right-hand side of Eq. (2) implies the following necessary condition: $\mathbf{f}_{x}\circ\mathbf{g}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{f}_{t}\circ\mathbf{g}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$ for all real pairs $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$. In the second step, utilizing the necessary condition and the conditional distribution defined in Eq. (4), we demonstrate that $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, defined as $\mathbf{h}$, and they are independent of the variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, which confirms that Eq. (7) and Eq. (8) do not depend on modality-specific latent variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. Finally, the third step concludes our result by introducing the inference model Eqs. (7)-(8). See Appendix A.4.1 for details.
Identifiability Discussion
Theorem 3.1 builds upon and extends the findings of Theorem 1 as presented in Zimmermann et al. (2021) to multi-modal context. Specifically, within a single-modal context, Theorem 3.1 aligns with the result outlined in Theorem 1 of Zimmermann et al. (2021). This advancement facilitates the application of Propositions 1 and 2 from Zimmermann et al. (2021) to multimodal contrastive representation learning. As a result, we can demonstrate that the minimization of Equation (6) in Theorem 3.1 is capable of identifying the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ up to a linear transformation, i.e., the recovered latent variable $\mathbf{\hat{z}}_{x}$, obtained through the minimization of Equation (6), is linearly related to the true $\mathbf{z}_{x}$ as follows: $\mathbf{\hat{z}}_{x}=\mathbf{A}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{A}$ represents an orthogonal matrix, and $\mathbf{c}$ is a constant vector. The comprehensive proof supporting this statement can be found in Appendix for completeness A.4.2.
The identifiability result above enables the extraction of independent latent variables through linear ICA after acquiring the recovered latent variables $\mathbf{\hat{z}}$. It is essential to acknowledge the geometric constraints imposed by the hypersphere setting. While it may not be feasible for all components of $\mathbf{z}$ to be independent, a subset of $\mathbf{z}$ can still exhibit independence. The extraction of these independent subsets is significant for investigating the capability of disentanglement. Further details on this subject will be thoroughly explored in Section 4.
@@ 3.2 Identifiability Analysis on convex bodies
The previous theoretical result requires the latent coupled variables to be a hypersphere, this will limit the disentangled ability of multimodal contrastive representation learning, due to the nature of the geometric constraints on the hypersphere. In this section, we will show a similar result for convex bodies, e.g., the hyperrectangle $[a_{1},b_{1}]\times...\times[a_{M},b_{M}]$, which allows for independence among the latent coupled variables, offering a more flexible framework.
On convex bodies, we parameterize the proposed latent partial causal generative models depicted in Figure 1 by the following:
$p(\mathbf{z}_{x})=|\mathcal{Z}_{c}|^{-1},\quad p(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{p}(\mathbf{z}_{x})e^{-\delta(\mathbf{z}_{t},\mathbf{z}_{x})/\lambda},$ (9)
$\mathbf{x}=\mathbf{g}_{x}(\mathbf{z}_{x},\mathbf{m}_{x}),\quad\mathbf{t}=\mathbf{g}_{t}(\mathbf{z}_{t},\mathbf{m}_{t}),$ (10)
where $\delta$ is a distance metric induced by a norm (See Appendix A.5.2). Diverging from the hypersphere space mentioned above, here we consider a convex body in $\mathbb{R}^{M}$, denoted as $\mathcal{Z}_{c}$. In this context, we assume that $p(\mathbf{z}_{x})$ follows a uniform distribution, and the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ follows an exponential distribution as defined in Eq. (9). Again, we do not enforce any further assumptions for $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. For $\mathbf{g}_{x}$ and $\mathbf{g}_{t}$, we assume them to be invertible mapping, ensuring information in latent space can be recovered. Given these assumptions on a convex body, similar to Theorem 3.1 on hypersphere, we have the following result:
@@
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eq. (9)-(10) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized symmetric contrastive loss given by Eq. (2) leads to the following reduction of the loss itself:
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$
$\quad\quad\quad+\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}$ (11)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},$ (12)
$q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},$ (13)
@@
Proof sketch
Similar to Proof sketch for theorem 3.1. Firstly, we demonstrate that by minimizing Eq. (2), the following necessary condition holds: $\mathbf{f}_{x}\circ\mathbf{g}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{f}_{t}\circ\mathbf{g}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$ holds for all pairs $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$. Next, based on the condition, we show that $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, defined as $\mathbf{h}$, and they are independent of the variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, which enables Eq. (12) and Eq. (13) not depending on modality-specific latent variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. Finally, the third step concludes our proof by introducing the model Eqs. (12)-(13). See Appendix A.5.1 for details.
Identifiability Discussion
Theorem 3.2 extends the findings of Theorem 5 in Zimmermann et al. (2021). Again, in the context of a single-modal setting, Theorem 3.1 is consistent with the results of Theorem 5 as stated in Zimmermann et al. (2021). Given this, we can easily demonstrate that Eq. (11) in theorem 3.2 can identify the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ up to a permutation transformation, i.e., the recovered latent variable $\mathbf{\hat{z}}_{x}$, obtained through the minimization of Equation (6), is linearly related to the true $\mathbf{z}_{x}$ as follows: $\mathbf{\hat{z}}_{x}=\mathbf{P}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{P}$ represents an orthogonal matrix, and $\mathbf{c}$ is a constant vector, by using Theorem 6 in Zimmermann et al. (2021). For completeness, the detailed proof of this assertion is provided in Appendix A.5.2.
4 Learning Disentangled Representations
Building on the foundation laid by our identifiability results, we now shift our focus to investigating the disentanglement ability by multimodal contrastive representation learning. Recognizing that training a model from scratch could be time-consuming and resource-intensive, we turn our attention to the disentanglement potential of the pre-trained multimodal model, e.g., CLIP. In this section, we propose two distinct methods, guided by our identifiability results, for exploring the disentanglement ability of the pre-trained CLIP model. These two methods are tailored to suit the different latent space settings of hypersphere space and convex bodies.
On hypersphere
Theorem 3.1 suggests that minimizing Eq. (6), in alignment with the standard training process of the CLIP model, enables the identification of latent variables $\mathbf{z}_{x}$ up to linear transformations. This is significant as it facilitates the extraction of independent latent variables from features acquired via a pre-trained CLIP model, further processed using linear ICA methods. An important consideration is the geometric constraints inherent to the hypersphere, particularly for a unit ${M-1}$ hypersphere. In this context, the maximum number of independent dimensions that can define each point on the hypersphere is $M-1$. Consequently, the latent variables $\mathbf{z}_{x}$ are limited to having at most $M-1$ independent dimensions. In the extreme situation, all true latent variables $\mathbf{z}_{x}$ might be dependent. Even in such scenarios, linear ICA methods may be still effective, since it is designed to extract components or features with maximum independence, utilizing the principle of maximal non-gaussianity (Hyvärinen et al., 2001). In our implementation, we employ the FastICA method proposed in Hyvarinen (1999).
On convex bodies
Unlike a hypersphere space, a convex body provides the flexibility to have independent latent variables $\mathbf{z}_{x}$. Our Theorem 3.2 demonstrates that, by minimizing Eq. (6) within a convex body setting, we can identify these latent variables $\mathbf{z}_{x}$ up to permutation. However, there exists a discrepancy: the CLIP model is originally trained in a hypersphere space, which inherently differs from the convex body assumptions foundational to Theorem 3.2. Despite this divergence, the pre-trained CLIP model may still be effective, since it, essentially, still aligns correct pairs and distances incorrect ones. It is important to note that the essential factor in achieving permutation identifiability through the minimization of Eq. (6) is the property of isometry in the mapping $\mathbf{h}$ in Theorem 3.2. Although a global isometric mapping from a convex body to a whole hypersphere is not feasible, the existence of a local and approximate isometry from a convex body to a small region on the hypersphere is a plausible consideration. Taking this into account, we can utilize Principal Component Analysis (PCA) to process the features learned by the pre-trained CLIP model, effectively reducing redundant information on the hypersphere. Subsequently, FastICA can be applied to counteract the orthogonal transformation effect introduced by PCA, thereby extracting the final features.
5 Experiments
Experiments on Synthetic Data
In our initial experiments, we use synthetic data to verify our main identifiability results on hypersphere and convex bodies, and empirically demonstrate the robustness of these results when facing substantial violations of assumptions. The synthetic data is generated through a specific process as the following. We consider latent coupled variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$, each with a dimensionality of 10. Additionally, we have modality-specific latent variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, both set to a dimension of 5. The process begins with sampling from the marginal distribution $p(\mathbf{z}_{x})$. We explore this both under conditions that align with our theoretical assumptions (using uniform distributions) and under conditions that deviate from these assumptions (using non-uniform distributions). We then create real pairs by sampling from the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$. This is done in two different settings: one that matches our assumptions about the conditional distribution and one that violates these assumptions. The observational data $\mathbf{x}$ and $\mathbf{t}$ are generated using two different 3-layer Multi-Layer Perceptrons (MLPs) with leaky ReLU units. Both sets of modality-specific latent variables, $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, are independently sampled from a Gaussian distribution. Beyond the hypersphere space, our experiments also encompass the bounded box (a type of convex body) and unbounded space.

### v2

@@ 4.1 Identifiability Analysis on Hypersphere
On hypersphere, we parameterize the proposed latent partial causal generative models as following:
$p(\mathbf{z}_{x})=|\mathcal{Z}|^{-1},\quad p(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{p}^{-1}e^{(k\mathbf{z}^{T}_{t}\mathbf{z}_{x})},\quad\mathbf{x}=\mathbf{g}_{x}(\mathbf{z}_{x},\mathbf{m}_{x}),\quad\mathbf{t}=\mathbf{g}_{t}(\mathbf{z}_{t},\mathbf{m}_{t}),$ (4)
where $\mathcal{Z}$ denotes the space of latent factors $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$. We assume that $\mathcal{Z}$ is the unit hypersphere $\mathbb{S}^{M-1}$, aligning with the commonly used normalization in constrastive loss. We do not enforce any further assumptions for $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. For $\mathbf{g}_{x}$ and $\mathbf{g}_{t}$, we assume them to be nonlinear, and invertible (i.e., injective) mapping, ensuring the information in latent space can be recovered. In addition, we assume that $p(\mathbf{z}_{x})$ follows a uniform distribution, and $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ follows a von Mises-Fisher (vMF) distribution, considering the constraint of unit hypersphere. Given these assumptions, we first establish that the minimization of the symmetric cross-entropy Eq. (2)) converges to a symmetric cross entropy, as follows:
@@
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eqs. (4) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized multimodal contrastiveloss given by Eq. (2) leads to the following reduction of the loss itself:
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N+2\log|\mathcal{Z}|=$ $\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x}),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t}),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}},$ (5)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})^{-1}e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)}q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})^{-1}e^{(\mathbf{h}(\mathbf{z}_{t})^{T}\mathbf{h}(\mathbf{z}_{x})/\tau)},$ (6)
@@
$C_{q}(\mathbf{z}_{t})=\int e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)}\mathrm{d}\mathbf{z}_{x},C_{q}(\mathbf{z}_{x})=\int e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)}\mathrm{d}\mathbf{z}_{t}.$
Refer to Appendix D.1 for proof. This is a generalization of Theorem 1 in [67].
Bridge Between Modalities
By addressing key asymmetries arising from modality differences, such as modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, along with distinct generative processes $\mathbf{g}_{x}$ and $\mathbf{g}_{t}$, we derive the result in Theorem 4.1. This result is pivotal as it establishes a critical connection between multimodal contrastive learning and traditional single-modal contrastive learning. In particular, Theorem 4.1 enables the transfer of insights and results from single-modal settings to the multimodal context. As a result, we present the following corollary:
Corollary 4.2.
By leveraging Theorem 4.1, the minimization of Eq. (5) identifies the latent variables $\mathbf{z}_{x}$ (and symmetrically, $\mathbf{z}_{t}$) up to a linear transformation. Specifically, the representations $\mathbf{f}_{x}(\mathbf{x})$, learned by the minimization of Eq. (5), are linearly related to the underlying latent variables $\mathbf{z}_{x}$ in the proposed latent partial causal model, as follows: $\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{A}$ is an orthogonal matrix and $\mathbf{c}$ is a constant vector.
For further details, see Appendix D.2.
Success of Multimodal Contrastive Learning
Corollary 4.2 shows that minimizing Eq. (5) (or equivalently, the multimodal contrastive loss in Eq. (1)) identifies the latent variables $\mathbf{z}_{x}$ (and symmetrically, $\mathbf{z}_{t}$) up to a linear transformation. This means that the representations $\mathbf{f}_{x}(\mathbf{x})$, learned through multimodal contrastive learning, are directly related to the latent variables $\mathbf{z}_{x}$ via a linear transformation, i.e., $\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}\mathbf{z}_{x}+\mathbf{c}$. A similar result holds for $\mathbf{z}_{t}$. This finding highlights the effectiveness of multimodal contrastive learning, suggesting that its success in practical applications stems from its ability to recover latent coupled variables. This recovery preserves essential, transferable knowledge across modalities, enabling the learned representations to capture high-level transferable information while discarding model-specific details. Such properties are key to the robustness and transferability of multimodal contrastive learning representations.
@@ 4.2 Identifiability Analysis on Convex Bodies
We now extend the previous identifiability result to convex bodies, e.g., the hyperrectangle $[a_{1},b_{1}]\times...\times[a_{M},b_{M}]$. On convex bodies, we parameterize the proposed generative models by the following:
$p(\mathbf{z}_{x})=|\mathcal{Z}_{c}|^{-1},\quad p(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{p}(\mathbf{z}_{x})^{-1}e^{-\delta(\mathbf{z}_{t},\mathbf{z}_{x})/\lambda},\quad\mathbf{x}=\mathbf{g}_{x}(\mathbf{z}_{x},\mathbf{m}_{x}),\quad\mathbf{t}=\mathbf{g}_{t}(\mathbf{z}_{t},\mathbf{m}_{t}),$ (7)
where $\delta$ is a distance metric induced by a norm. We consider a convex body in $\mathbb{R}^{M}$, denoted as $\mathcal{Z}_{c}$, where we assume that $p(\mathbf{z}_{x})$ follows a uniform distribution, and the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ follows an exponential distribution. Again, we do not enforce any further assumptions for $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. For $\mathbf{g}_{x}$ and $\mathbf{g}_{t}$, we assume them to be nonlinear and invertible mapping, ensuring information in latent space can be recovered. Given these assumptions, we have the following result:
@@
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eq. (7) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized multimodal contrastiveloss given by Eq. (2) leads to the following reduction of the loss itself:
$\small\lim_{N\rightarrow\infty}\mathcal{L}-2\log N+2\log|\mathcal{Z}_{c}|=\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x}),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t}),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}},$ (8)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},$ (9)
@@
$C_{q}(\mathbf{z}_{t})=\int e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau}\mathrm{d}\mathbf{z}_{x},C_{q}(\mathbf{z}_{x})=\int e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau}\mathrm{d}{\mathbf{z}_{t}}.$
Bridge Between Modalities
In convex bodies, Theorem 4.3, introduced for the first time in this work, plays a key role in bridging multimodal contrastive learning with traditional contrastive learning by addressing the asymmetric challenges arising from modality differences. Building on this theorem, we can seamlessly extend the findings of Theorem 5 from the single-modal context [67] to the multimodal setting, as follows:
Corollary 4.4.
By leveraging Theorem 4.3, the minimization of Eq. (8) in theorem 4.3 identifies the latent variables $\mathbf{z}_{x}$ (symmetrically, $\mathbf{z}_{t}$) up to a permutation transformation, i.e., the representations $\mathbf{f}_{x}(\mathbf{x})$, learned by the minimization of Eq. (8), is related to the underlying $\mathbf{z}_{x}$ in the proposed partial causal model as follows: $\mathbf{f}_{x}(\mathbf{x})=\mathbf{P}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{P}$ is an permutation matrix with scaling, $\mathbf{c}$ is a constant vector.
For completeness, see details in Appendix E.2.
Success of Multimodal Contrastive Learning
Corollary 4.2 Similar to Corollary 4.2 on hyperspheres, Corollary 4.4 establishes that, on convex bodies, the representations $\mathbf{f}_{x}(\mathbf{x})$ learned by multimodal contrastive learning are related to the true latent variables $\mathbf{z}_{x}$ as $\mathbf{f}_{x}(\mathbf{x})=\mathbf{P}\mathbf{z}_{x}+\mathbf{c}$. This provides a theoretical foundation for the success of multimodal contrastive learning on convex bodies.
5 From Principles to Practice: Disentanglement Ability of CLIP-like Models
In theory, both Corollaries 4.2 and 4.4 suggest a disentanglement potential of CLIP-like models trained by multimodal contrastive learning, under the assumption that the variables in $\mathbf{z}_{x}$ (and symmetrically, $\mathbf{z}_{t}$) are mutually independent. we explore how these theoretical insights can be translated into practical guidance for the effective use of CLIP-like models.
Corollary 4.2 shows that the representations $\mathbf{f}_{x}(\mathbf{x})$ learned by multimodal contrastive learning are linearly related to the true latent variables $\mathbf{z}_{x}$ via an orthogonal transformation, i.e., $\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}\mathbf{z}_{x}+\mathbf{c}$. This result holds under two key conditions: (1) the true latent variables are sampled from a hyperspherical latent space, and (2) the inference model, e.g., CLIP-like models, is trained in a hyperspherical inference space. Notably, CLIP-like models naturally satisfy condition (2), as they typically employ L2 normalization, constraining representations to the unit sphere. Therefore, under condition (1) holds, the representations from CLIP-like models can be passed through linear unmixing method (e.g., FastICA [24]) to resolve the mixing matrix $\mathbf{A}$, resulting in disentangled representations. It is worth to note that the geometry of the hypersphere, specifically the unit ${M-1}$-dimensional hypersphere, places an upper bound on the number of independent variables, i.e., ${M-1}$ at most.
Unlike Corollary 4.2, Corollary 4.4 shows that the learned representations $\mathbf{f}_{x}(\mathbf{x})$ from multimodal contrastive learning are already disentangled, i.e., $\mathbf{f}_{x}(\mathbf{x})=\mathbf{P}\mathbf{z}_{x}+\mathbf{c}$. This result mainly requires two conditions: (1) the true latent variables $\mathbf{z}_{x}$ are sampled from a convex body latent space, and (2) the inference space of the model is also constrained to a convex body. However, CLIP-like models typically violate the second condition, as they operate in a hyperspherical inference space due to L2 normalization, even through we assume that the first condition may hold. Nevertheless, the insight from Corollary 4.4 remains useful with appropriate adjustments. In particular, the Corollary relies on the existence of an isometric mapping from the latent space to the representation space (see Eq. 48 in the Appendix). Although a global isometry from a convex body to a hypersphere is not feasible, it is reasonable to assume a local isometry between the convex body and small regions of the hypersphere. Based on this, we propose first applying Principal Component Analysis (PCA) to the representations $\mathbf{f}_{x}(\mathbf{x})$. Then, FastICA can be used to account for the orthogonal transformation introduced by PCA, enabling the extraction of the final disentangled representations. This PCA+ICA pipeline thus enables effective use of CLIP-like models under the result of Corollary 4.4.
By leveraging the disentanglement capabilities of CLIP-like models, we can improve performance on tasks that benefit from disentangled representations, such as few-shot learning and domain generalization. We hope this motivates further exploration of the disentanglement potential inherent in CLIP-like models across a broad range of downstream applications.
6 Synthetic Experiments and Real-World Evaluation with Pretrained CLIP
Synthetic Experiments
In our initial experiments, we use synthetic data to validate our main identifiability results on hyperspheres and convex bodies, while also empirically assessing their robustness under significant violations of assumptions. We first sample $p(\mathbf{z}_{x})$ according to the distributions listed in Table 3. Additionally, we generate paired samples from the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ following the distributions specified in the same table. Beyond hyperspheres, our experiments also consider bounded and unbounded spaces. Each experiment is repeated three times for every setting. For more details regarding experiments, refer to Appendix K.